OpenStudy (princesssleelee):

Given that f(x) = −x + 4 and g(x) = −2x − 3, solve for f(g(x)) when x = 2. A. −14 B.−7 C. 7 D. 11

OpenStudy (princesssleelee):

I'm thinking 11

OpenStudy (solomonzelman):

yes, correct.

OpenStudy (princesssleelee):

Thank you :)

OpenStudy (solomonzelman):

Example Problem: \(\large\color{#000000 }{ \displaystyle f(x)=3x-5 }\) \(\large\color{#000000 }{ \displaystyle g(x)=5x+4 }\) Find f(g(x)) when x=3. SOLUTION: your function f(x) is \(\large\color{#000000 }{ \displaystyle f(\color{red}{x} )=3\color{red}{x}-5 }\) to find f( g(x) ) you need to plug the g(x) into the f(x), for x. \(\large\color{#000000 }{ \displaystyle f(\color{red}{g(x)} )=3(\color{red}{g(x)})-5 }\) But, I know that g(x) is 5x+4, so I can apply this for my right side. \(\large\color{#000000 }{ \displaystyle f(\color{red}{g(x)} )=3(\color{red}{5x+4})-5 }\) \(\large\color{#000000 }{ \displaystyle f(\color{red}{g(x)} )=15x+12-5 }\) \(\large\color{#000000 }{ \displaystyle f(\color{red}{g(x)} )=15x+7 }\) and when x=3, we get \(\large\color{#000000 }{ \displaystyle f(\color{red}{g(3)} )=15(3)+7=45+7=52 }\)